OpenStudy (anonymous):

Give the range of x - 2y = 0 for the domain {-1, 0, 4}.

OpenStudy (mertsj):

The range is the set of y values. You have been given x values so plug each x into the given equation and find the y value. Those 3 y values will be the range for the domain (x values) given.

OpenStudy (amistre64):

the range is the set of values that are a result of operations done with the dependant variables

OpenStudy (amistre64):

as is, there is not indication as to if y is a function of x; or x is a function of y, or if there is some other function of x,y

OpenStudy (amistre64):

*done with independant .... :)

OpenStudy (mertsj):

So if x = -1 then -1 -2y = 0 and y = -1/2 If x = 0, y = 0 If x = 4 then 4-2y = 0 and -2y = -4 and y = 2